in terms of pi, how much farther does a 70cm diameter bicycle tire travel in one revolution than a 50cm diameter tire
Answer by jim_thompson5910(35256)

in terms of pi, how much farther does a 70cm diameter bicycle tire travel in one revolution than a 50cm diameter tire


Circumference of the 70cm diameter tire

C = pi*r^2

C = pi*(35)^2 ... note: if the diameter is 70 cm, then the radius is 70/2 = 35 cm

C = pi*1225

C = 1225pi

---------------------------------------------------------------------------------------

Circumference of the 50cm diameter tire

C = pi*r^2

C = pi*(25)^2 ... note: if the diameter is 50 cm, then the radius is 50/2 = 25 cm

C = pi*625

C = 625pi

---------------------------------------------------------------------------------------

Now subtract the two circumferences (large - small): 1225pi - 625pi = (1225 - 625)pi = 600pi

The difference is 600pi, which is the final answer.